OpenStudy (anonymous):

Write the logarithm in expanded form. (Enter the logarithmic function in the form log_a(b).) log7(5y)

OpenStudy (anonymous):

\[\log(5)+\log(y)\]

OpenStudy (anonymous):

where did the 7 go?

OpenStudy (anonymous):

oh the seven is the base and i didn't write it because it is not relevant to the problem, but your answer is \[\log_7(5)+\log_7(y)\]

OpenStudy (anonymous):

\[\log_b(xy)=\log_b(x)+\log_b(y)\] irrespective of the base

OpenStudy (anonymous):

\[\frac{\text{Log}[5]+\text{Log}[y]}{\text{Log}[7]} \]

OpenStudy (anonymous):

Log is to the base e in the expression above.